Octane (C3H18) is a major component of gasoline. When octane burns in excess oxygen water is produced as well as carbon dioxide. The density of gasoline is 0.79gmL-1. What volume of carbon dioxide(at STP) is produced from the combustion of 1.0L of octane.
1
Expert's answer
2020-10-23T09:15:54-0400

The equation of the given reaction is

2C8H18 + 25O2 --> 16CO2 + 18H2O


The mass of octane equals "0.79g/mL\times1000mL=790g"


"n (CO_2)=\frac{790g (C_8H_{18})}{114.2g/mol}\times\frac{16mol(CO_2)}{2mol(C_8H_{18})}=55.3mol"


"V(CO_2)=55.3mol\times22.4L/mol=1.24\cdot10^3L"


Answer: "1.24\cdot10^3L"